P3-2A Nona Curry started her own consulting firm, Curry Consulting Inc., on May 1, 2017. The following transactions occurred during the month of May.1 Stockholders invested S15,000 cash in the business in exchange for common stock. 2 Paid $600 for office rent for the month. 3 Purchased $500 of supplies on account. 5 Paid $150 to advertise in the County News 9 Received $1,400 cash for services performed. 12 Paid $200 cash dividend. 15 Performed $4,200 of services on account. 17 Paid $2,500 for employee salaries. May 20 Paid for the supplies purchased on account on May 3 Received a cash payment of $1,200 for services performed on account on May 15. Borrowed S5,000 from the bank on a note payable. Purchased office equipment for $2,000 paying $200 in cash and the balance on account. Paid $180 for utilities. 23 26 29 30

Please analyze transactions with a assets=liabilities+stockholders' equity table. Ending cash balance equals $18,270, while total assets come out to $23,770. Net income totals to $2,170
Answer #1
Date Assets = Liabilities + Stockholders' Equity
Cash Accounts Receivable Supplies Equipment = Notes Payable + Accounts Payable Common Stock + Retained Earnings
Revenue - Expenses - Dividends
May 1 $15,000 + + + = + + $15,000 + - -
2 -$600 + + + = + + + - $600 -
3 + + $500 + = + $500 + + - -
5 -$150 + + + = + + + - $150 -
9 $1,400 + + + = + + + $1,400 - -
12 -$200 + + + = + + + - - $200
15 + $4,200 + + = + + + $4,200 - -
17 -$2,500 + + + = + + + - $2,500 -
20 -$500 + + + = + -$500 + + - -
23 $1,200 + -$1,200 + + = + + + - -
26 $5,000 + + + = $5,000 + + + - -
29 -$200 + + + $2,000 = + $1,800 + + - -
30 -$180 + + + = + + + - $180 -
$18,270 + $3,000 + $500 + $2,000 = $5,000 + $1,800 + $15,000 + $5,600 - $3,430 - $200
Net income $2,170